Entry for Factory Labor Costs

The weekly time tickets indicate the following distribution of labor hours for three direct labor employees:

Hours
Job 301 Job 302 Job 303 Process
Improvement
Tom Couro 19 10 5 5
David Clancy 9 15 11 4
Jose Cano 10 14 9 5

The direct labor rate earned per hour by the three employees is as follows:

Tom Couro $20
David Clancy 22
Jose Cano 18

The process improvement category includes training, quality improvement, and other indirect tasks.

a. Journalize the entry to record the factory labor costs for the week. If an amount box does not require an entry, leave it blank.
Answer #1

a.

To record entry for factory labor costs is:

Journal Entry 13 Date Particulars Debit Credit 14 Work in Process A/c Factory Overhead A/c To Wages Payable A/c (To record wa

Working Notes:

a.

Computation of total number of hours of T is:

Total number of hours of T = Number of hours for Job 301 + Number of hours for Job 302 + Number of hours for Job 303

= 19 + 10 + 5

= 34

Hence, the total number of hours of T is 34.

b.

Computation of total number of hours of D is:

Total number of hours of D = Number of hours for Job 301 + Number of hours for Job 302 + Number of hours for Job 303

= 9 + 15 + 11

= 35

Hence, the total number of hours of D is 35.

c.

Computation of total number of hours of J is:

Total number of hours of J = Number of hours for Job 301 + Number of hours for Job 302 + Number of hours for Job 303

= 10 + 14 + 9

= 33

Hence, the total number of hours of J is 33.

d.

Computation of work in process is:

Work in process = (Total number of hours of T * Direct labor rate of T) + (Total number of hours of D * Direct labor rate of D) + (Total number of hours of J * Direct labor rate of J)

= (34 * $20) + (35 * $22) + (33 * $18)

= $680 + $770 + $594

= $2,044

Hence, the work in process is $2,044.

e.

Computation of indirect labor is:

Indirect labor = (Process improvement hours of T * Direct labor rate of T) + (Process improvement hours of D * Direct labor rate of D) + (Process improvement hours of J * Direct labor rate of J)

= (5 * $20) + (4 * $22) + (5 * $18)

= $100 + $88 + $90

= $278

Hence, the indirect labor is $278.